Ticket Number: HEA-NOC/20090318-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: Airspeed
Ticket Opened: 20090318 09:40 UTC Problem Start: 20090317 19:00 UTC
Ticket Update: 20090401 09:41 UTC+1 Problem End: 20090401 08:33 UTC+1
Intermittent connectivity to GMIT-Castlebar
Intermittent connectivity to GMIT with 2 min outages between flaps
20090318 09:31 COC Contacted Airspeed for an alarm check.
20090318 13:02 COC Airspeed have identified the problem which will require
an on-site investigation and a possible further disruption of services.
Clients will be contacted before any maintenance is carried out
20090330 09:47 LT Further issues arose with this link over the weekend,
with outages of a few minutes at a time occurring from 17h04 on Saturday
until the link dropped at 04h17 this morning and has remained down since.
Airspeed engineers are at their site in Sraheen at the moment, working to
resolve the problem.
20090330 10:36 GAR Airspeed confirm that the fault was due to water
gaining ingress to the cabling. This may be related to a previous problem
with water gaining ingress to the ODU recently, and a question has been
raised to the manufacturers re correct installation procedure to ensure
that this does not happen again.
The cable has been replaced and the high-site equipment is being
recalibrated at the moment and is expected back online shortly.
20090330 12:29 LT Airspeed have completed their works and the link has now
been restored. We will continue to monitor for 24 hours.
20090331 09:43 LT The link is down once more. It started to flap at 21h38
yesterday, with outages of a couple of minutes at a time occuring several
times this morning.
Airspeed suspect the cause is remaining water in the IF cable in Sraheen
and have sent engineers to the site to swap the cable. They expect them to
be on site in 3 hours. Their first priority is to restore connectivity,
after which they will build a parallel link and will swap over to the new
link once done. The swapover will incur another brief outage but they
believe that the new link will resolve all of the problems at the site.
This issue is currently being escalated internally with a view to ensuring
that the root cause is correctly identified and resolved once and for all.
20090331 13:43 LT Airspeed have been at the Sraheen site for the last 40
minutes or so but have been unable to restore the link. They are now
proceeding with running a new IF cable to implement a new link - they
expect it to take approximately one hour to commission this new link.
20090331 16:10 LT The link is back up since 15h59. We are awaiting
feedback from Airspeed to confirm whether repair works are now complete and
whether they believe the link is now stable.
20090331 17:59 LT The link remains unstable. Airspeed will replace the
antenna at Sraheen tomorrow morning. The work is expected to take about 1
hour and will commence at approximately 09h00. Airspeed are confident that
this will resolve the persistent problems with the link.
20090401 09:38 LT Airspeed arranged for the engineer to get to site
earlier than originally planned (at 07h20) and they completed the work and
brought the link back up ahead of schedule at 08h33. The link has remained
stable and error free since then, according to both our monitoring system
and that of Airspeed, but we will continue to monitor it closely over the
next few days.
This ticket can be monitored at http://www.hea.net/tickets/20090318-1
HEAnet Limited. Registered in Ireland, No. 275301.